The mass of aluminium liberated by 1 faraday
1
Expert's answer
2020-03-09T09:42:23-0400

The equation of aluminum electrolysis is


"\text{Al}^{3+}+3\bar{e}\rightarrow \text{Al}^0."

Therefore, 1 mole of aluminum requires 3 moles of electrons. Keep in mind that 1 faraday is not a SI unit and equals 96485.33 Coulombs, however, this unit is easier to use. According to the reaction, in the equation below "z=3":


"m=F\frac{M}{z}=1\cdot\frac{27\text{ g}}{3}=9\text{ g}."